OpenStudy (anonymous):

-10 + x ÷ 2(9-x^2) for x = -6

OpenStudy (anonymous):

Is there a shorter way to do this problem?

OpenStudy (anonymous):

\[\frac{-10 + x }{2\times (9 - x^2} = \frac{-10-6}{2 \times (9 -(-6)^2)} = \frac{-16}{2 \times (9-36)}= \frac{16}{54}\]

OpenStudy (anonymous):

\[\frac{16}{54} = \frac{8}{27}\]

OpenStudy (anonymous):

Hey, thanks Sheg. :):)

OpenStudy (anonymous):

its ok buddy :)

OpenStudy (anonymous):

\[-10 + x/2 (9 - x^2) \]\[-10+\frac{1}{2} x \left(9-x^2\right) \]Replace x with -6.\[-10+\frac{1}{2} (-6) \left(9-(-6)^2\right) \]\[-10+(-3) (-27) \]\[71 \] http://www.wolframalpha.com/input/?i=+-10+%2B+x%2F2+%289+-+x%5E2%29%2F.+x-%3E-6